25 June 2000  Page 1 of 1  J3/00-235

Subject: Issues 270 and 271 From: Van Snyder

1 Introduction

Issues 270 and 271 are both concerned with type aliases. Issue 270 asks whether changing declaration-type-spec to type-spec is the correct change to prevent creating aliases for classes. The answer is “yes.” Issue 271 asks whether something should be said about the impossibility of using a type alias as a parent type name in an extended type definition. The answer is “yes.”

2 Edits

Edits refer to 00-007r2. Page and line numbers are displayed in the margin. Absent other instructions, a page and line number or line number range implies all of the indicated text is to be replaced by immediately following text, while a page and line number followed by + indicates that immediately following text is to be inserted after the indicated line. Remarks for the editor are noted in the margin, or appear between [ and ] in the text.

[Editor: Replace “name” with “type-name” to emphasize that an alias can’t be used here.] 41:21

[Editor: Remove “declaration-”.] 57:19

[Editor: Delete issues 270 and 271.] 57:24-33

NOTE 4.49\frac{1}{2} 57:35+ It is not possible to use a type alias name as the parent type name in the declaration of an extended type (4.5.3).